A 15-year annuity pays $2,000 per month, and payments are made at the end of each month. The interest rate is 11 percent compounded monthly for the first Six years and 9 percent compounded monthly thereafter.

Required:

What is the present value of the annuity?

  1. $181,632.49
  2. $185,265.14
  3. $2,179,589.93
  4. $177,999.84
  5. $252,753.46 <This was wrong
Answer #1

first calculate the value after 6 years

FV = 0, PMT = 2000, rate = 9%/12, N = 9*12

use PV function in Excel

value after 6 years = 147,678.7632

FV = 147,678.7632, N = 6*12, PMT = 2000, rate = 11%/12

use PV function in Excel

present value = 181,632.49